I have a lemonade stand and charged $1.00 per cup and sold 48 cups. When I raised the price to $2.00 per cup I only sold 27 cups. I need to write an equation for the number of cups I sold as a function of the price I charged. "C" would be for the number of cups and "P" would denote the price I charged. I have to assume the function is linear. I am stumped ---- can you help???
Answer by stanbon(75887)

I have a lemonade stand and charged $1.00 per cup and sold 48 cups. When I raised the price to $2.00 per cup I only sold 27 cups. I need to write an equation for the number of cups I sold as a function of the price I charged. "C" would be for the number of cups and "P" would denote the price I charged.
----
You have two points (1,48) and (2,27)
----
slope = (48-27)/(1-2) = -21
-----------------------
Form: C = mP + b
Substitute and solve for "b":
48 = -21(1)+b
b = 69
----
Equation:
C(P) = -21P + 69
======================
Cheers,
Stan H.
=======================